Water under the ice
Yesterday, Steff and Trevor went down with the HT ice core drill and found rock. Today they went down with the rock drill and found … ice. It appears that the deepest part of the borehole is filling up with water that freezes in.
In just two days enough ice has formed that a hollow ice core came up today with the rock drill. The ice forms from the sides of the bore hole despite a drill liquid overpressure. Whereas this is an interesting finding in itself, it may put an end to the rock/sediment drilling, because the hole freezes in faster that it can be drilled. In two weeks time, when the rock drilling is supposed to continue the deepest part of the bore hole is likely to be full of ice.
Weather conditions have been improving today. The wind, snowfall, and snow drift has disappeared and instead we got sunshine and warm temperatures.
The good weather inspired a lot of camp activity: Lou and Sverrir have been on the skiway with the beam and tiller, Lars went out to do a few more GPS measurements, Eliza dug a 2m tephra pit, Astrid, Christine, and Nanna checked and repaired the skiway flags, Bruno packed the last electronics boxes, and for the first time Motohiro did not have to dig his way into the yellow vapour sampling tent.
As there is absolutely no wind and the sun is out the dome rapidly warms up, so despite open doors and windows the temperature on the third floor became close to 30 °C. Outside the temperature went up to -3 °C, but with the disappearing clouds we have good chance of a cool night.
What we have done today:
1. Went down with rock drill and found ice
2. Skiway and apron grooming with beam and tiller
3. NEGIS: yesterday’s weather was a bit rough but today, the sun is back
and it should be a productive day
4. Oil change on main generator
5. Checking skiway flags
6. Clearing entrances to trenches and tents
7. GPS work around camp
8. 2 m tephra pit study
Ad. 2: It was snowing and drifting all night, so in the morning the skiway was full of fresh, soft snow. Lou started beam grooming at 10 AM after the snowing had stopped. Sverrir went out with the tiller at around noon. In the evening the skiway conditions are good. With a cold night it should be in good condition for receiving the skier.
Ad. 3: The oil was changed from 8:20 to 8:40 in the morning. In the meantime the vapor sampling system in the yellow tent was powered from a small generator
In the morning, frequent snow showers, drifting snow and wind up to 10 knots from W. During the day the wind dropped to <5 knots from S, the temperature went up to -3 °C, and we had sunny spells/broken clouds. At midnight we have -9 °C and ground fog.